# Forecasting Treatment Response with Deep Pharmacokinetic Encoders

### v7

@@ C.4 Forecasting with Dose-Independent Parameters
Baseline Exog. Baseline Exog.
@@
C.5 Model MAE and RMSE Results
This appendix presents the complete results for all models for both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis. Table 6 provides complete results for the simulated data and Table 7 provides complete results for the OhioT1DM dataset.
Table 6: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exog. Baseline Exog.
MAE
ETS 9.450 13.457
(0.00) (0.00)
MLP 10.589 8.433 14.699 10.600
@@ C.5 Model MAE and RMSE Results
RNN 19.033 27.116 28.834 43.256
@@
NHITS-ST 10.082 12.679
(0.054) (0.135)
NHITS-PK 10.039 12.339
(0.188) (0.435)
Table 7: Mean absolute error (MAE) (top) and Root Mean Square Error (RMSE) (bottom) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exog. Baseline Exog.
MAE
ETS 9.022 10.028
(0.00) (0.00)
MLP 11.065 11.120 12.630 12.698
@@
LSTM 9.811 9.467 11.252 11.178
(0.570) (0.230) (0.512) (0.396)
TCN 11.525 10.760 13.647 12.678
(2.008) (0.925) (3.349) (1.459)
NHITS 9.060 8.873 10.337 10.137
(0.142) (0.064) (0.137) (0.108)
NHITS-ST 8.958 10.131
(0.087 ) (0.100)
NHITS-PK 8.758 9.965
(0.080) (0.095)
RMSE
ETS 14.569 16.225
(0.00) (0.00)
MLP 16.534 16.948 18.651 19.289
(1.280) (0.481) (1.403) (0.943)
RNN 16.498 17.066 19.446 19.423
(1.252) (2.037) (1.709) (2.532)
LSTM 14.915 14.597 16.894 16.851
(0.494) (0.279) (0.490) (0.394)
TCN 16.690 15.971 19.213 18.418
(1.891) (0.710) (2.706) (1.295)
NHITS 14.844 14.505 16.789 16.453
(0.192) (0.096) (0.189) (0.132)
NHITS-ST 14.640 16.492
(0.205) (0.203)
NHITS-PK 14.385 16.251
(0.092) (0.089)

### v8

@@ C.2 Forecasting with Dose-Independent Parameters
w/o Exog. w/ Exog. w/o Exog. w/ Exog.
@@
[…]
We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values. However, this approach led to worse performance on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters (Xie, 2018; Visentin et al., 2016). In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3. Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data. These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
C.3 Model MAE and RMSE Results
This appendix presents the complete results for all models for both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis. Tables 8 and 9 provide complete MAE and RMSE results, respectively.
Table 8: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset. Average results are computed across 8 trials. Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based, and PK model groups, respectively. Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable. Best results are highlighted in bold, second best results are underlined. PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
Simulated Dataset OhioT1DM Dataset
Model All Values Critical Values All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
AutoETS 9.450 13.457 9.022 10.028
@@ C.3 Model MAE and RMSE Results
Table 9: Root Mean Square Error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset. Average results are computed across 8 trials. Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based groups, and PK model groups, respectively. Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable. Best results are highlighted in bold, second best results are underlined. PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
Simulated Dataset OhioT1DM Dataset
Model All Values Critical Values All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
AutoETS 14.751 21.454 14.569 16.225
@@
NBEATSx-PK 9.480 11.244 16.945 17.659